OpenStudy (anonymous):

Is there a bijection \(f:\mathbb R^+\to\mathbb R^+\) s.t. \(f'(x)=f^{-1}(x)\)?

OpenStudy (anonymous):

ill help

OpenStudy (anonymous):

Existence can be shown because \(f(x)=\phi^{-\frac\phi{\phi+1}}x^\phi\) satisfies the conditions. Is this unique?

OpenStudy (anonymous):

\[f'(f(x))=x,\forall x\in I:=(0,\infty),\therefore f\in C^\infty(I)\] Is \(f\) analytic?

OpenStudy (anonymous):

Never mind, I was just on my way to M.SE I guess.

OpenStudy (anonymous):

Oh wow, check this out. http://en.wikipedia.org/wiki/Bernstein%27s_theorem_on_monotone_functions

OpenStudy (anonymous):

Never mind, I suppose I'm just talking to myself here...

OpenStudy (anonymous):

THANKS FOR ALL THE HELP. Just kidding, I love you.

OpenStudy (anonymous):

lol